OpenStudy (anonymous):

Suppose a parabola has an axis of symmetry at x = –7 , a maximum height of 4 and also passes through the point (–6, 0). Write the equation of the parabola in vertex form.

DivineSolar (divinesolar):

Please post in the Mathematics section not English. @PrincessBee

OpenStudy (anonymous):

Oh haha!! Oops sorry! Kind of new to this!

DivineSolar (divinesolar):

@PrincessBee Anyways i will help you with this one but anyways here is the answer. The vertex form of the equation is y = a(x - h)2 + k Since the vertex is defined as (h, k), h = -7 and k = 4: y = a(x - (-7))2 + 4 or: y = a(x + 7)2 + 4 We just need to find the value of "a" (which must be negative, remember, it opens down). Here is where we use the point through which the parabola passes through: (-6, 0). We substitute the values x = -6, y = 0 in the equation and solve for a: 0 = a(-6 + 7)2 + 4 simplify and subtract 4 on both sides of the equation: -4 = a(1)2 or a = -4 Now we are ready to write our completed equation: y = -4(x + 7)2 + 4 Medal if this helped.

OpenStudy (anonymous):

So y=-4(x+7)2+4 is the answer!?

DivineSolar (divinesolar):

Yes.

DivineSolar (divinesolar):

Medal? Did i help :D?

DivineSolar (divinesolar):

If you dont know how to do that then you click best response to the right of my name. @PrincessBee